OpenStudy (anonymous):

Choose the correct answer that shows the base 10 number 23 in binary form.

OpenStudy (perl):

there are algorithms online, but 23 is between two powers of 2 , 2^5 < 23 <= 2^4

OpenStudy (anonymous):

i dont get it @perl

OpenStudy (perl):

a binary number is basically powers of 2 a1 *2^n + a2*2^n-2 + ... an , where an is either 0 or 1

OpenStudy (mathmath333):

10=1010 10+10=1010+1010=10100 10+10+3=10111

OpenStudy (mathmath333):

http://www.mathsisfun.com/binary-digits.html

OpenStudy (aum):

2 | 23 ------ 2 | 11 R: 1 ------ 2 | 5 R: 1 ------ 2 | 2 R: 1 ------ 2 | 1 R: 0 ------ 1 R: 1 R is remainder at each stage of division. Write the remainders in a sequence from the last remainder to the first one: 10111

OpenStudy (mathmath333):

oh thats nice method i forgot

OpenStudy (perl):

repeated division , and noting the remainders

OpenStudy (perl):

upside down division is a nice touch